# DESY/PUBDB-2024-01826

### v1

@@ 3.1 V0 Vertex Reconstruction
The vertex finding algorithm employed in this study first identifies and removes all the tracks originating from the primary vertex or V0 candidates.
The V0s are found by reconstructing all possible vertices with pairs of oppositely charged tracks and constraining their invariant masses, displacements, and directions. The vertices are not discarded but stored and assigned a particle ID based on the set of constraints that they pass, summarised in Table 1. Three processes are considered: $K_{S}^{0}\to\pi^{+}\pi^{-}$, $\Lambda^{0}\to p\pi^{-}$, and photon conversions, $\gamma_{conv}$ $X\to e^{+}e^{-}X$. The invariant mass of these reconstructed V0s can be seen in Figure 1. The mass of the decay particles used to calculate the invariant mass of the V0 is decided based on which set of constraints the V0 passes, unlike the SV, where all tracks are assumed to be pions in the invariant mass calculation.
$K_{S}^{0}$ $\Lambda^{0}$ $\gamma_{conv}$
tight loose tight loose tight loose
M [GeV] $[0.493,0.503]$ $[0.488,0.508]$ $[1.111,1.121]$ $[1.106,1.126]$ $<0.005$ $<0.01$
r [mm] $>0.5$ $>0.3$ $>0.5$ $>0.3$ $>9$ $>9$
$\hat{\text{p}}\cdot\hat{\text{r}}$ $>0.999$ $>0.999$ $>0.99995$ $>0.999$ $>0.99995$ $>0.999$
Table 1: Summary of the default V0 selection criteria Suehara_2016 . M is the invariant mass, and p is the momentum of the V0 candidate. r is the distance of the V0 candidate from the primary vertex. $\hat{\text{p}}\cdot\hat{\text{r}}$ is defined as a collinearity variable for the V0 candidate. The set of ‘tight’ constraints has been used to identify V0s in this study, while the set of ‘loose’ constraints has been used to reconstruct the seed vertex.
@@ 3.2 Secondary Vertex Reconstruction
Secondary Vertices are found by reconstructing a two-track vertex (seed) with the lowest $\chi^{2}$ from the vertex fit and attaching tracks to this seed, resulting in a vertex with the lowest $\chi^{2}$ until the resulting vertex no longer passes the criteria mentioned in Ref. Suehara_2016 . The tracks forming the SV are stored and removed from the original set, and more SVs are reconstructed recursively until no more seeds pass the required constraint thresholds.
Due to the near-diagonal CKM matrix, the cascading decay chain of heavier quarks is expected to be $b\to c\to s\to(u,d)$. Hence, the SV multiplicity tends to be higher in $b$-jets compared to $c$- and light jets, as can be seen in Figure 2(a). The resolution of the flight distance of the SV achieved using this reconstruction in $B_{S}^{0}$ decays can be seen in Figure 2(b). The flight distance resolution is defined as the difference between the radial distance of the reconstructed SV from the primary vertex and the radial distance of the MC decay vertex. The closest SV is associated with the MC decay vertex for events with multiple reconstructed SVs.
(a) SV multiplicity in jets from $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$ events
(b) Flight distance resolution in $10^{4}$ $B_{S}^{0}\rightarrow J/\psi\phi\rightarrow\mu\mu KK$ events.
Figure 2: Performance of secondary vertex reconstruction in processes generated at $\sqrt{s}=91.2$ GeV.
@@ 4 DeepJetTransformer
Since the introduction of ParticleNet, the concept of Particle Cloud has become the prevailing representation of jet structure. A Particle Cloud considers the jet as an unordered set of jet constituents of varying length. Elements of differing nature, such as charged, neutral particles, or SVs associated with the jet, are considered to create the most complete and accurate representation. This representation concept was used to build the presented model, the key element of which, the unordered set of particles, requires the construction of a model invariant under the permutation of the jet constituents. This essential property is in opposition to most Transformer models established around the principle of causality https://doi.org/10.48550/arxiv.1706.03762 ; devlin2019bert . It was also expected to design a model capable of extracting connections between the jet constituents, enhancing its capabilities of constructing engineered high-level features by capturing dependencies inside the global structure of the jet.
A structure based on Transformer blocks was thus chosen for this study. Previous research has indicated that Transformer models offer enhanced performance and increased efficiency, particularly compared to Graph models Qu:2022mxj . The subsequent sections will elaborate on the inputs to the neural network and the fundamental characteristics of Transformer models and provide a detailed description of the specific model, DeepJetTransformer, which has been developed for this study.
@@ 4.1 Input Features
The properties of each jet and its constituents represent different categories of input features available for model training. The jet kinematics are represented by variables defined using its $4-$momentum, as detailed in Table 2. Many future collider detector concepts are designed to be used with a particle flow algorithm BUSKULIC1995481 ; CMS-PAS-PFT-09-001 . Therefore, jet constituents are subdivided into five sets according to the typical particle flow candidate categories: charged hadrons, neutral hadrons, electrons and positrons $(e^{\pm})$, photons $(\gamma)$, and muons $(\mu^{\pm})$. Kinematic variables are defined for each jet constituent using its $4$-momentum, as listed in Table 3. Charged tracks are first fitted to find the V0s and the remaining tracks are used to reconstruct SVs. Feature variables are defined separately for both classes of reconstructed vertices (V0s and SVs) and are listed in Table 4. The distinguishing power of some of these variables is discussed below.
@@
$|p|,~{}E,~{}m$ momentum, energy, and invariant mass of the jet
@@
jet angularity Larkoski:2014pca as sum of normalized jet contituent energy ($z_{i}$)
$\lambda^{\kappa}_{\beta}=\Sigma_{i\in\text{jet}}z^{\kappa}_{i}\Theta^{\beta}_{i}$ and angular distance to jet axis ($\Theta_{i}$) for ($\kappa=0$, $\beta=0$),
@@
$\Theta$ angular distance of track and jet axis
@@
$\dfrac{|p|_{\text{ch.}}}{|p|_{\text{jet}}},\ln(|p|_{\text{ch.}}),\ln\left(\dfrac{|p|_{\text{ch.}}}{p|_{\text{jet}}}\right)$ (normalised) magnitude of track momentum and logarithms
@@
$m_{\text{neut.}}$ invariant mass of neutral particle
@@
$\Theta$ angular distance of neutral particle and jet axis
$\dfrac{|p|_{\text{neut.}}}{|p|_{\text{jet}}},\ln(|p|_{\text{neut.}}),\ln\left(\dfrac{|p|_{\text{neut.}}}{p|_{\text{jet}}}\right)$ (normalised) magnitude of particle momentum and logarithms
@@
$p,~{}m$ momentum and invariant mass of the SV
@@
Figure 3: Distinguishing features in the clustered jets of $e^{+}e^{-}\to Z\to q\bar{q}$ events at $\sqrt{s}=91.2$ GeV, separated by flavour. Fig 3(b) corresponds to the jet constituents, the rest correspond to the clustered jets. The IDEA detector concept was used for reconstruction.
The jet momentum distribution of $b$-jets tends to be more spread out than that of light jets, as seen in Figure 3(a). This is due to the longer decay chain in $b$-jets, where more momentum can be lost through neutrinos than in light jets.
An important distinguishing variable for $b$-jet identification is the transverse impact parameter ($D_{0}$), which is higher for heavier flavour jets as the decaying $B$ hadrons have a significantly longer lifetime than $D$ or light hadrons (except for V0s). The differentiating effect between flavours caused by this can be seen more clearly in the transverse impact parameter significance, defined as $S(D_{0})=D_{0}/\sigma_{D_{0}}$, where $\sigma_{D_{0}}$ is the uncertainty in the measurement of the transverse impact parameter. It is depicted in Figure 3(b).
As mentioned in Section 3.2, $b$-jets tend to have a higher SV multiplicity than $c$- and light jets. It is a dominant property in identifying $b$-jets and, to some extent, $c$-jets.
The most challenging background for $s$-tagging is $ud$-jets. Two powerful distinguishing variables tend to be the multiplicities of charged and neutral Kaons and Pions, exploiting the conservation of strangeness during hadronisation in strange jets. These can be seen in Figure 3(c) and 3(d). To distinguish between $K^{\pm}$ and $\pi^{\pm}$, PID techniques like energy loss ($dE/dx$) Lippmann2012 , ionisation cluster counting ($dN/dx$) Cataldi:1996mz , time-of-flight Klempt:1999tx , etc. are traditionally used. The $K^{\pm}$/$\pi^{\pm}$ classification is performed with a varying efficiency of correctly identifying $K^{\pm}$, the baseline scenario being $90\%$ efficiency, and a $10\%$ efficiency of misidentifying $\pi^{\pm}$ as $K^{\pm}$. The baseline PID scenario was deliberately conservative with respect to the state-of-the-art $K^{\pm}$ identification and follows PID studies at Belle, which found the average efficiency and fake rate for charged particles between $0.5$ and $4$ GeV/$c$ to be $(87.99\pm 0.12)\%$ and $(8.53\pm 0.10)\%$, respectively NAKANO2002402 . To further improve PID, neutral strange hadrons ($K_{S}^{0}$ and $\Lambda^{0}$) are reconstructed in the form of V0s, as was shown in Figure 1. These variables, as described in Table 2, 3, and 4, are fed into a neural network, the architecture of which is described below.
@@ 4.2 Transformer Models
Inspired by the success of attention mechanism in Natural Language Processing (NLP) https://doi.org/10.48550/arxiv.1706.03762 ; devlin2019bert or Computer Vision (CV) dosovitskiy2021image tasks, this model adopts Transformer blocks as its primary architectural component. Transformers belong to a class of neural networks that leverage the scaled dot-product attention mechanism https://doi.org/10.48550/arxiv.1706.03762 . The attention mechanism enables the model to selectively focus on specific segments of the input sequence while processing each constituent element. In contrast to earlier architectures, such as recurrent models that utilise fixed-size windows or recurrent connections, the attention mechanism dynamically assigns weights to individual elements within the jet based on their relevance, capturing intricate dependencies across the entirety of the jet structure. This adaptive and global weighting scheme empowers the Transformer to effectively model contextual information, a crucial element for understanding and generating coherent high-level features.
@@ 4.2.1 Scaled Dot-Product Attention and Heavy Flavour Transformer Block
The scaled dot-product attention (SDPA) mechanism uses three inputs: a query matrix $Q$, a key matrix $K$, and a value matrix $V$. The query matrix represents the items for which the attention weights are computed, while the key and value matrices represent all items in the sequence. After being fed into linear layers, the query tensor $Q$ of dimension $(B,N,d_{k})$, the key tensor $K$ of dimension $(B,L,d_{k})$, and the value tensor $V$ of dimension $(B,L,d_{k}^{\prime})$ are fed into the scaled dot-product attention as:
@@
The attention mechanism in this study is employed in a specific configuration where the query, key, and value tensors are identical. This particular case is commonly referred to as self-attention. The output tensors $Q$, $K$, and $V$ are generated through linear layers, facilitating the transformation and projection of the input tensors to the attention space.
SDPA is extended to enhance the discriminating power of the model by allowing it to attend to multiple subspaces of attention in parallel. This extension, referred to as Multi-Head Attention (MHA), facilitates the capture of diverse and complementary high-level features from the input by projecting the Query, Key, and Value matrices independently for each of the $h$ attention heads. Each attention head performs an SDPA operation, yielding distinct representations. These head representations are then concatenated and passed through a linear layer to integrate the information across heads. The MHA layer can mathematically be represented by the following equations:
@@
The presented approach, employing the Particle Cloud representation, intentionally refrains from employing positional encoding. This decision stems from the absence of a hierarchical structure or positional ordering among the components of the jets, in contrast to sequences such as sentences or image patches. Consequently, the MHA module operates without incorporating positional encoding and instead only leverages permutation invariant mechanisms to capture and process the interrelationships between particles in the jet, yielding meaningful results. By analogy with graph structures, the attention mechanism can be interpreted similarly to the ones used in fully connected Graph Networks, with the attention scores playing a role similar to the edge features by capturing relationships within the jet structure.
@@
  • The inputs are fed into a basic Multilayer Perceptron (MLP) layer followed by a ReLU activation function.
@@
  • In addition to the MHA layer, a fully connected feed-forward layer is also added, identical to the original Transformer implementation https://doi.org/10.48550/arxiv.1706.03762 followed by a last residual connection and layer normalisation.
Unlike other Transformer models applied to jet (sub)structures, a $cls$ token is not employed to embed the information of the jet structures into relevant features for classification. Instead, an attention pooling allowing is introduced, behaving similarly to a Max or Average pooling layer with an attention mechanism and learnable parameters. The attention pooling operates by employing an MLP projection layer, which enables local feature extraction. Subsequently, a softmax activation function is applied to calculate attention weights, allowing the layer to emphasise relevant elements in the sequence. The attention weights are then used to aggregate the sequence information by performing a weighted sum. To enhance the layer’s performance, batch normalisation is applied to normalise, the ReLU 4082265 activation function is used to introduce non-linearity, and dropout regularisation is incorporated to prevent overfitting. The attention pooling layer can effectively capture essential information from the sequence and produce a condensed representation by incorporating these components that can be utilised for jet flavour classification. DeepJetTransformer could also be interpreted as a fully connected graph network using the jet’s constituents as the nodes and the SDPA as a mechanism connecting all the node information for enhancing the feature engineering of the model.

### v4

@@ 3.1 V0 Vertex Reconstruction
Two processes are considered: $K_{S}^{0}\to\pi^{+}\pi^{-}$ and $\Lambda^{0}\to p\pi^{-}$. The invariant mass of the reconstructed $K_{S}^{0}$s can be seen in Figure 1(a), demonstrating a good reconstruction of V0s and their properties. The mass of the tracks used to calculate the invariant mass of the V0 is decided based on the set of constraints the V0 passes with a certain permutation of the two tracks. In contrast, all tracks are assumed to be pions in the invariant mass calculation for the SVs.
$K_{S}^{0}$ $\Lambda^{0}$
tight loose tight loose
M [GeV] $[0.493,0.503]$ $[0.488,0.508]$ $[1.111,1.121]$ $[1.106,1.126]$
r [mm] $>0.5$ $>0.3$ $>0.5$ $>0.3$
$\hat{\text{p}}\cdot\hat{\text{r}}$ $>0.999$ $>0.999$ $>0.99995$ $>0.999$
Table 1: Summary of the default V0 selection criteria Suehara_2016 . M is the invariant mass, and p is the momentum of the V0 candidate. r is the distance of the V0 candidate from the primary vertex. The collinearity of the V0 candidate is defined as $\hat{\text{p}}\cdot\hat{\text{r}}$. The set of ‘tight’ constraints has been used to identify V0s in this study, while the set of ‘loose’ constraints has been used to remove the V0 background while reconstructing SVs.
Figure 1(b) displays the V0 multiplicity in jets from $Z\to q\bar{q}$ events. No reconstructed V0s are found for most of the jets. But, a higher fraction of heavy- and strange-flavoured jets contain reconstructed V0s than $u$- and $d$-jets, which justifies the importance of V0 rejection before attempting to reconstruct SVs. It is also evident that more $s$-jets have one or more reconstructed V0s than $u$- and $d$-jets, making V0s an important discriminator of $s$-jets against lighter quark jets.
@@ 3.2 Secondary Vertex Reconstruction
Due to the near-diagonal CKM matrix, the cascading decay chain of heavier quarks is expected to be $b\to c\to s\to(u,d)$. Hence, the SV multiplicity tends to be higher in $b$-jets compared to $c$-, $s$-, $u$-, and $d$-jets, as shown in Figure 2.
Figure 2: SV multiplicity in jets from $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$ events. The term “light jets” here collectively refers to $u$-, $d$-, and $s$-jets.
@@ 4 DeepJetTransformer
Since the introduction of ParticleNet Qu:2019gqs , the concept of a Particle Cloud has become the prevailing representation of jet structure. A Particle Cloud considers the jet as an unordered set of jet constituents of varying length. Elements of differing nature, such as charged, neutral particles, or SVs associated with the jet, are considered to create the most complete and accurate representation. This representation concept was used to build the presented model, the key element of which, the unordered set of particles, requires the construction of a model invariant under the permutation of the jet constituents [^1] Permutation invariance is in opposition to most Transformer models established around the principle of causality vaswani2023attentionneed ; devlin2019bert ., a benefit also used by other transformer-based taggers Qu:2022mxj . Moreover, Transformers possess the essential property of full connectivity between jet constituents via the attention mechanism vaswani2023attentionneed . This enables the model to capture subtle correlations among jet constituents, enhancing the high-level features used for jet discrimination.
A structure based on Transformer blocks was thus chosen for this study. Previous research has indicated that Transformer models offer enhanced performance and increased efficiency, particularly compared to graph models Qu:2022mxj ; 10363595 . The subsequent sections will elaborate on the inputs to the neural network and the fundamental characteristics of Transformer models and provide a detailed description of the specific model, DeepJetTransformer, which has been developed for this study.
@@ 4.1 Input Features
The properties of each jet and its constituents represent different categories of input features available for model training. All input features are built using information reconstructed with Delphes detector simulation unless stated otherwise. The jet kinematics are represented by variables defined using its $4-$momentum, as detailed in Table 2. Many future collider detector concepts are designed to be used with a particle flow algorithm BUSKULIC1995481 ; CMS-PAS-PFT-09-001 . Therefore, jet constituents are subdivided into five sets according to the typical particle flow candidate categories: charged hadrons, neutral hadrons, electrons and positrons $(e^{\pm})$, photons $(\gamma)$, and muons $(\mu^{\pm})$. Kinematic variables are defined for each jet constituent using its $4$-momentum, as listed in Table 3. For each jet up to 25 charged jet constituents and 25 neutral jet constituents are considered. This is enforced by truncating the input feature array of a given jet if the number of charged/neutral jet constituents is more than 25. Conversely, if the number of charged/neutral jet constituents is less than 25, then the input feature array is zero-padded.
Charged tracks are first fitted to find the V0s and the remaining tracks are used to reconstruct SVs. Feature variables are defined separately for both classes of reconstructed vertices (V0s and SVs) and are listed in Table 4. Up to 4 V0s and 4 SVs are considered per jet. The V0 and SV input feature arrays are likewise truncated/zero-padded. The distinguishing power of some of these variables is discussed below.
@@
$|p|,~{}E,~{}m$ 3-momentum magnitude, energy, and invariant mass of the jet
@@
jet angularity Larkoski:2014pca as sum of normalized jet constituent energy ($z_{i}$)
$\lambda^{\kappa}_{\beta}=\Sigma_{i\in\text{jet}}z^{\kappa}_{i}R^{\beta}_{i}$ and angular distance to jet axis ($R_{i}$) for ($\kappa=0$, $\beta=0$),
@@
$R$ angular distance of track and jet axis
@@
$\dfrac{|p|_{\text{ch.}}}{|p|_{\text{jet}}},\ln(|p|_{\text{ch.}}),\ln\left(\dfrac{|p|_{\text{ch.}}}{|p|_{\text{jet}}}\right)$ (normalised) magnitude of track momentum and logarithms
@@
$R$ angular distance of neutral particle and jet axis
$\dfrac{|p|_{\text{neut.}}}{|p|_{\text{jet}}},\ln(|p|_{\text{neut.}}),\ln\left(\dfrac{|p|_{\text{neut.}}}{|p|_{\text{jet}}}\right)$ (normalised) magnitude of particle momentum and logarithms
@@
$|p|,~{}m$ 3-momentum magnitude and invariant mass of the SV
@@
Figure 3: Distinguishing features in the clustered jets of $e^{+}e^{-}\to Z\to q\bar{q}$ events at $\sqrt{s}=91.2$ GeV, separated by flavour. Fig 3(b) shows a property of the jet constituents, while the rest show properties of the clustered jets. The IDEA detector concept was used for reconstruction.
The jet 3-momentum magnitude distribution of $b$- and $c$-jets tends to be more spread out than that of $s$-, $u$-, and $d$-jets, as seen in Figure 3(a). This is due to the longer decay chain in $c$-jets than $s$-, $u$-, and $d$-jets, and even longer decay chains in $b$-jets, where more momentum can be lost through neutrinos than in $s$-, $u$-, and $d$-jets.
An important distinguishing variable for $b$-jet identification is the transverse impact parameter ($D_{0}$), which is higher for heavier flavour jets as the decaying $b$ hadrons have a significantly longer lifetime than $c$ or $s$, $u$, $d$ hadrons (except for V0s). The differentiating effect between flavours caused by this can be seen more clearly in the transverse impact parameter significance, defined as $S(D_{0})=D_{0}/\sigma_{D_{0}}$, where $\sigma_{D_{0}}$ is the uncertainty in the measurement of the transverse impact parameter. It is depicted in Figure 3(b).
As mentioned in Section 3.2, $b$-jets tend to have a higher SV multiplicity than $c$-, $s$-, $u$-, and $d$-jets. It is a dominant property in identifying $b$-jets and, to some extent, $c$-jets.
The most challenging background for $s$-tagging is $ud$-jets. Two powerful distinguishing variables tend to be the multiplicities of charged and neutral Kaons and Pions, exploiting the conservation of strangeness during hadronisation in strange jets. These can be seen in Figure 3(c) and 3(d). To distinguish between $K^{\pm}$ and $\pi^{\pm}$, PID techniques like energy loss ($dE/dx$) Lippmann2012 , ionisation cluster counting ($dN/dx$) Cataldi:1996mz , time-of-flight Klempt:1999tx , etc. are traditionally used. The $K^{\pm}$/$\pi^{\pm}$ classification is generically emulated, instead of relying on any particular PID technique, with several scenarios of different efficiency to correctly identify $K^{\pm}$, the baseline scenario being $90\%$ efficiency and a $10\%$ efficiency of misidentifying $\pi^{\pm}$ as $K^{\pm}$. The $K^{\pm}$ identification efficiency and the $\pi^{\pm}$ misidentification efficiency are chosen to be constant over the entire momentum range for all the scenarios. The other scenarios considered to study the impact of PID on flavour tagging are summarised in Section 5.2. The baseline PID scenario was deliberately conservative with respect to the state-of-the-art $K^{\pm}$ identification, which is expected to provide better than $3\sigma$ $K^{\pm}$/$\pi^{\pm}$ separation using cluster counting at FCC-ee CAPUTO2023167969 , potentially supplemented by time-of-flight Benedikt:2651299 ; Wilkinson2021 ). This study instead follows PID studies at Belle, which found the average efficiency and fake rate for charged particles between $0.5$ and $4$ GeV/$c$ to be $(87.99\pm 0.12)\%$ and $(8.53\pm 0.10)\%$, respectively NAKANO2002402 . The reconstructed V0s, as shown in Figure 1, further improve PID by identifying the neutral strange hadrons, $K_{S}^{0}$ and $\Lambda^{0}$. These variables, as described in Table 2, 3, and 4, are fed into a neural network, the architecture of which is described below.
@@ 4.2 Transformer Models
Inspired by the success of attention mechanism in Natural Language Processing (NLP) vaswani2023attentionneed ; devlin2019bert or Computer Vision (CV) dosovitskiy2021image tasks, this model adopts Transformer blocks as its primary architectural component. Transformers belong to a class of neural networks that leverage the scaled dot-product attention (SDPA) mechanism vaswani2023attentionneed . The attention mechanism enables the model to selectively focus on specific segments of the input sequence while processing each constituent element. In contrast to earlier architectures, such as recurrent models that utilise fixed-size windows or recurrent connections, the attention mechanism dynamically assigns weights to individual elements within the jet based on their relevance, capturing intricate dependencies across the entirety of the jet structure. This adaptive and global weighting scheme empowers the Transformer to effectively model contextual information, a crucial element for understanding and generating coherent high-level features.
@@ 4.2.1 Scaled Dot-Product Attention and Heavy Flavour Transformer Block
The SDPA mechanism uses three inputs: a query matrix $Q$, a key matrix $K$, and a value matrix $V$. In general, the query matrix represents the items for which the attention weights are computed, while the key and value matrices represent all items in the sequence. In this study, the items can be understood to be jet constituents. After being fed into linear layers, the query tensor $Q$ of dimension $(B,N,d_{k})$, the key tensor $K$ of dimension $(B,L,d_{k})$, and the value tensor $V$ of dimension $(B,L,d_{k}^{\prime})$ are fed into the scaled dot-product attention as:
@@
The attention mechanism in this study is employed in a specific configuration where the input query, key, and value tensors are identical ($Q=K=V$), and derived from jet constituent features. The tensors Q, K, and V are each passed through linear layers, facilitating the transformation and projection of the input tensors to the attention space. The SDPA is then computed on these transformed tensors as $\text{Attention}(QW^{Q},KW^{K},VW^{V})$, where $W^{Q}$, $W^{K}$, $W^{V}$, represent distinct linear transformations. This particular case is commonly referred to as self-attention vaswani2023attentionneed .
SDPA is extended to enhance the discriminating power of the model by allowing it to attend to multiple subspaces of attention in parallel. This extension, referred to as Multi-Head Attention (MHA), facilitates the capture of diverse and complementary high-level features from the jet constituent input by projecting the Query, Key, and Value matrices independently for each of the $h$ attention heads. Each attention head performs an SDPA operation, yielding distinct representations. These head representations are then concatenated and passed through a linear layer to integrate the information across heads. The MHA layer can mathematically be represented by the following equations:
@@
The presented approach, employing the Particle Cloud representation Qu:2019gqs , intentionally refrains from employing positional encoding. This decision stems from the absence of a hierarchical structure or positional ordering among the components of the jets, in contrast to sequences such as sentences or image patches. Consequently, the MHA module operates without incorporating positional encoding and instead only leverages permutation invariant mechanisms to capture and process the interrelationships between particles in the jet, yielding meaningful results. The permutation invariance of DeepJetTransformer is established by the properties of permutation equivariance and invariance of function composition xu2024permutationequivariancetransformersapplications . The permutation equivariance of each function of the transformer blocks ensures that the network produces a representation of the jet constituents respecting the Particle Cloud properties. It is made sure that the network’s flavour predictions remain invariant under the permutation of jet constituents by applying a permutation invariant attention pooling followed by linear layers for classification. By analogy with graph structures, the attention mechanism can be interpreted similarly to the ones used in fully connected graph networks, with the attention scores playing a role similar to the edge features by capturing relationships within the jet structure.
@@
  • The jet constituent inputs are fed into a basic Multilayer Perceptron (MLP) layer followed by a ReLU activation function.
@@
  • In addition to the MHA layer, a fully connected feed-forward layer is also added, identical to the original Transformer implementation vaswani2023attentionneed followed by a last residual connection and layer normalisation.
Unlike other Transformer models applied to jet (sub)structures Mondal_2024 ; Qu:2022mxj , a $cls$ token is not employed to embed the information of the jet structures into relevant features for classification. Instead, an attention pooling is introduced, behaving similarly to a Max or Average pooling layer with an attention mechanism and learnable parameters. The attention pooling operates by employing an MLP projection layer, which enables local feature extraction. Subsequently, a softmax activation function is applied to calculate attention weights, allowing the layer to emphasise relevant elements in the sequence. The attention weights are then used to aggregate the sequence information by performing a weighted sum. To enhance the layer’s performance, batch normalisation is applied, the ReLU 4082265 activation function is used to introduce non-linearity, and dropout regularisation is incorporated to prevent overfitting. The attention pooling layer can effectively capture essential information from the sequence and produce a condensed representation by incorporating these components that can be utilised for jet flavour classification. In the context of jet flavour tagging, Transformer models can be interpreted as fully connected graph networks using the jet’s constituents as the nodes, and the SDPA as a mechanism connecting all the node information for enhancing the feature engineering of the model.